# Misaligned incentives in the post-2024 UEFA Champions League: A managerial failure

### v5

@@ 5 Conclusions
Our paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. The order in which the vacancies caused by the titleholders are filled and the two European performance spots are allocated has almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt one of our recommendations in order to safeguard the integrity of its flagship tournament.
@@
Date/Period Issue
@@
May 2012 The match Feyenoord vs Heerenveen is played, where Heerenveen was better off by losing than by playing a draw (Csató,, 2021, Chapter 2.1)
@@
2015/16–2023/24 Incentive incompatibility of the UEFA Champions League seeding (Csató, 2020b, )
@@
The problem identified is especially worrying because it is rooted in the allocation of vacant slots, a topic that was repeatedly responsible for incentive incompatibility in UEFA club competitions as summarised in Table 2—while the fair solution has already been provided by Dagaev and Sonin, (2018). This issue offers an instructive lesson for decision-makers around the world: commitment to sporting merit requires not only media releases but appropriately designed tournaments. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented. In order to achieve this fundamental aim, Kendall and Lenten, (2017) suggest two alternative calls to action:
@@ Acknowledgements
One anonymous colleague provided valuable comments and suggestions on an earlier draft.

### v6

@@ 6 Conclusions
Our paper has revealed that the UEFA Champions League rules are incentive incompatible again from the 2024/25 season. In particular, the allocation rules of the vacancies caused by the titleholders and the two European performance spots have almost led to a game with misaligned incentives in the 2023/24 German Bundesliga. UEFA is strongly encouraged to adopt our recommendation in order to safeguard the integrity of its flagship tournament.
@@
Date/Period Problem
@@
May 2012 The match Feyenoord vs. Heerenveen, where Heerenveen was better off by losing than by playing a draw (Csató,, 2021, Chapter 2.1)
@@
2015/16–2023/24 Incentive incompatibility of the UEFA Champions League seeding system (Csató, 2020b, )
@@
The problem identified is especially worrying because it is rooted in the allocation of vacant slots, a topic that was repeatedly responsible for incentive incompatibility in UEFA club competitions as summarised in Table 2. This issue offers an instructive lesson for decision-makers around the world: commitment to sporting merit requires not only media releases but appropriately designed tournaments. Since rule changes such as the introduction of the European performance spot can always affect fairness, a thorough assessment is necessary before they are implemented.
In order to achieve that fundamental aim, Kendall and Lenten, (2017) suggest two alternative calls to action:
@@ Acknowledgements
Phil Scarf and one anonymous colleague provided valuable comments and suggestions on an earlier draft.